# Hybrid Reinforcement:When Reward Is Sparse, It’s Better to Be Dense

### v1

@@ A.1 Experimental setup
Actor Model Base model 1 OctoThinker-8B
@@
Table 7: Key hyperparameters used for GRPO training on OpenMathReasoning (Moshkov et al., 2025) in the verl (Sheng et al., 2025) framework for the OctoThinker-8B.
@@
For hybrid reward training for both Qwen-4B-Base and OctoThinker-8B-Base, we set the range parameters $\alpha$ and $\beta$ depending on the task type. For verifiable tasks, we adopt a tighter setting $\alpha=\beta=0.05$ to exploit the high precision of rule-based verifiers while minimizing noise. For mixed and hard-to-verify tasks, where the reward model contributes more substantially to supervision, we relax the range to $\alpha=\beta=0.1$ to provide richer feedback. For variance-aware reweighting, we fix the weighting bounds as $w_{\min}=0.4$ and $w_{\max}=3.0$, with a steepness parameter $k=6$ in the logistic weighting function. These values ensure that trivial prompts are down-weighted, while highly uncertain prompts—where reward-model scores vary widely—receive stronger emphasis without destabilizing training.
@@
Tables 6 and 7 provide an overview of the hyperparameter configurations used in our GRPO training runs with Qwen3-4B-Base and OctoThinker-8B. The tables cover settings across data preparation, actor model optimization, rollout generation, reward specification, and trainer configuration. They highlight the consistent use of OpenMathReasoning as the training corpus, the integration of both rule-based and reward-model signals, and the adoption of scalable rollout and training strategies within the verl framework. Together, these summaries document the experimental setup and ensure reproducibility across different backbone models. In addition, we employ the HuggingFace math_verify library to provide standardized rule-based verification of responses against ground-truth answers, which guarantees consistency in supervision across all experiments.
@@ A.3 More experiments
Table 8: Comparison with model-based verifiers on Qwen-3-4B-Base. Results are pass@1 averaged over 8 seeds across verifiable and hard-to-verify reasoning tasks. HERO consistently outperforms both General Reasoner and Qwen2.5-7B-IT under all regimes.
Verifiable tasks Hard-to-verify tasks
@@
Training with verifiable samples
@@
To further assess whether hybrid reward learning can outperform existing model-based verifiers, we compare HERO against two representative systems: General Reasoner, a frozen 1.5B verifier model that provides binary correctness judgments (Ma et al., 2025), and Qwen2.5-7B-IT, a large instruction-tuned verifier (Yang et al., 2024b). As shown in Table 2, HERO consistently achieves higher accuracy than both model-based verifiers under all training regimes. When trained with verifiable samples, HERO attains an average score of 62.0, outperforming General Reasoner (58.4) and Qwen2.5-7B-IT (62.5) while maintaining greater stability across datasets such as MATH500 (85.4 versus 82.8 and 83.7) and AMC (69.4 versus 62.8 and 58.1). In the hard-to-verify regime, the advantage becomes more pronounced: HERO reaches 56.5, exceeding General Reasoner (54.0) and Qwen2.5-7B-IT (54.5), demonstrating that hybrid reward learning provides more reliable supervision even when symbolic verification is unreliable and model-based signals are uncertain. In the mixed setting, which combines both verifiable and open-ended samples, HERO again leads with 64.1, surpassing General Reasoner (59.0) and Qwen2.5-7B-IT (63.3). These results highlight that integrating verifier-anchored and reward-model signals yields not only better accuracy but also more consistent generalization across regimes, outperforming larger model-based verifiers despite using no additional model parameters or external training data. The improvement underscores that structured reward integration, rather than sheer verifier scale, is the key to effective and robust reasoning optimization.
@@
Methods Verifiable Hard-to-verify

### v3

@@ A.1 Experimental setup
Actor Model Base model 1 OctoThinker-8B-Hybrid-Base
@@
Table 7: Key hyperparameters used for GRPO training on OpenMathReasoning (Moshkov et al., 2025) in the verl (Sheng et al., 2025) framework for the OctoThinker-8B-Hybrid-Base.
@@
For hybrid reward training for both Qwen-4B-Base and OctoThinker-8B-Hybrid-Base, we set the range parameters $\alpha$ and $\beta$ depending on the task type. For easy-to-verify tasks, we adopt a tighter setting $\alpha=\beta=0.05$ to exploit the high precision of rule-based verifiers while minimizing noise. For mixed and hard-to-verify tasks, where the reward model contributes more substantially to supervision, we relax the range to $\alpha=\beta=0.1$ to provide richer feedback. For variance-aware reweighting, we fix the weighting bounds as $w_{\min}=0.4$ and $w_{\max}=3.0$, with a steepness parameter $k=6$ in the logistic weighting function. These values ensure that trivial prompts are down-weighted, while highly uncertain prompts—where reward-model scores vary widely—receive stronger emphasis without destabilizing training.
@@
Tables 6 and 7 provide an overview of the hyperparameter configurations used in our GRPO training runs with Qwen3-4B-Base and OctoThinker-8B-Base. The tables cover settings across data preparation, actor model optimization, rollout generation, reward specification, and trainer configuration. They highlight the consistent use of OpenMathReasoning as the training corpus, the integration of both rule-based and reward-model signals, and the adoption of scalable rollout and training strategies within the verl framework. Together, these summaries document the experimental setup and ensure reproducibility across different backbone models. In addition, we employ the HuggingFace math_verify library to provide standardized rule-based verification of responses against ground-truth answers, which guarantees consistency in supervision across all experiments.
@@ A.3 More experiments
Table 8: Comparison with model-based verifiers on Qwen-3-4B-Base. Results are pass@1 averaged over 8 seeds across verifiable and hard-to-verify reasoning tasks. HERO consistently outperforms both General Reasoner and Qwen2.5-7B-Instruct under all regimes.
Easy-to-verify tasks Hard-to-verify tasks
@@
Training with easy-to-verify samples
@@
To further assess whether hybrid reward learning can outperform existing model-based verifiers, we compare HERO against two representative systems: General Reasoner, a frozen 1.5B verifier model that provides binary correctness judgments (Ma et al., 2025), and Qwen2.5-7B-Instruct , a large instruction-tuned verifier (Yang et al., 2024b). As shown in Table 2, HERO consistently achieves higher accuracy than both model-based verifiers under all training regimes. When trained with verifiable samples, HERO attains an average score of 62.0, outperforming General Reasoner (58.4) and Qwen2.5-7B-Instruct (62.5) while maintaining greater stability across datasets such as MATH500 (85.4 versus 82.8 and 83.7) and AMC (69.4 versus 62.8 and 58.1). In the hard-to-verify regime, the advantage becomes more pronounced: HERO reaches 56.5, exceeding General Reasoner (54.0) and Qwen2.5-7B-Instruct (54.5), demonstrating that hybrid reward learning provides more reliable supervision even when symbolic verification is unreliable and model-based signals are uncertain. In the mixed setting, which combines both verifiable and open-ended samples, HERO again leads with 64.1, surpassing General Reasoner (59.0) and Qwen2.5-7B-Instruct (63.3). These results highlight that integrating verifier-anchored and reward-model signals yields not only better accuracy but also more consistent generalization across regimes, outperforming larger model-based verifiers despite using no additional model parameters or external training data. The improvement underscores that structured reward integration, rather than sheer verifier scale, is the key to effective and robust reasoning optimization.
@@
Methods Easy-to-verify Hard-to-verify